OpenStudy (anonymous):

HELP!!! Twice the difference of a number and 9 equals 3 . Use the variable b for the unknown number.

OpenStudy (anonymous):

let number be=x thus differnce of number =x-9 twice difference= 2(x-9)=3

OpenStudy (anonymous):

so its 2(x-9) = 3

OpenStudy (anonymous):

Since they said use b, it will be 2(b-9)=3

OpenStudy (anonymous):

oh yeah b right sorry

OpenStudy (anonymous):

The sum of a number times 7 and 8 is equal to 5. Use the variable w for the unknown number.

OpenStudy (anonymous):

7w+8=5

OpenStudy (anonymous):

your really good cn u keep helpin

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Six less than the quotient of a number and 7 is equal to 2. Use the variable b for the unknown number.

OpenStudy (anonymous):

@mohammad.arqum

OpenStudy (anonymous):

(b/7)-6=2

OpenStudy (anonymous):

Four times the sum of a number and 9 is equal to 3 . Use the variable y for the unknown number.

OpenStudy (anonymous):

4(y+9)=3

OpenStudy (anonymous):

The difference of a number times 6 and 8 is equal to 3. Use the variable y for the unknown number.

OpenStudy (anonymous):

6y-8=3

OpenStudy (anonymous):

THANK U!! ;) my other ones are much harder dont think u could do but thanks u

OpenStudy (anonymous):

hmm